# Ultra-cold atoms as quantum simulators for relativistic phenomena

### v1

@@ 2.4.1   Analogue Universe in condensate at rest
Instead of a monotonically expanding Universe, an oscillating scale factor has been simulated in the Chicago experiment [145] by modulating the external magnetic field near a Feshbach resonance [103]. By measuring the number of excited atoms in situ, it was possible to observe the thermal character of the radiation emitted into a certain direction. In terms of the coupled harmonic oscillator picture sketched in Fig. 2.1, the reduced density matrix of one oscillator is thermal, i.e., a mixed state, even though the total quantum state of both oscillators is a pure (entangled) state. Indeed, the coherence associated to this two-mode squeezed state could also be tested in the Chicago experiment [145]. Note that the title of the publication [145] refers to Unruh radiation, which can probably be explained by the fact that squeezing underlies both phenomena, cosmological particle creation and the Unruh effect (as well as many others), but the analogy to cosmological particle creation as considered in this section is apparently closer and more quantitative.
@@ 2.4.2   Expanding condensates
An expanding ring-shaped Bose-Einstein condensate has been realized in the Maryland experiment [150]. Since the wave-length of the phonons expands with the condensate, such a set-up is well suited for observing the red-shift of the field modes within an expanding Universe. In the Maryland experiment [150], this has been achieved by imprinting a sound wave (i.e., a classical excitation) in the initial state of the condensate. As a related but more subtle phenomenon, the Maryland experiment [150] did also find evidence for Hubble friction. To explain this phenomenon, let us consider the equation of motion (2.1) for a mass-less scalar field in the Friedmann-Lemaitre-Robertson-Walker metric (2.18) in 3+1 dimensions in terms of the proper time $\tau$
@@
After the initial period of linear evolution, the Maryland experiment [150] was also able to observe the subsequent non-linear dynamics, which are qualitatively analogous to the pre-heating and re-heating periods in the early Universe after the end of inflation.
@@ 2.5   Gibbons-Hawking effect
at the expense of introducing coordinate singularities at the de Sitter horizon $r=1/H$. In 3+1 dimensions, we would find the same form with an additional $r^{2}d\Omega^{2}$ for the angular part.
@@ 2.7   Super-radiance and Penrose process
So far, the phenomenon of super-radiance discussed above refers to purely classical wave dynamics. Actually, it even has an analogue in terms of classical particles, which is the Penrose process [159] and allows us to extract energy from a rotating black hole. Furthermore, the phenomenon of super-radiance can be generalized to other scenarios such as a rotating body [160]. Generally speaking, super-radiance requires an effective energy which can become negative, some sort of scattering or tearing apart process and some mechanism to absorb the parts with negative energy. For surface waves [158], this phenomenon has actually been observed for such a vortex flow (2.28) in the Nottingham experiment [161]. Replacing the water as used in the Nottingham experiment [161] by a “photon super-fluid”, an analogous effect has been observed in the Glasgow experiment [162], see also [163] and [164].
@@ 2.7.2   Quasi-normal modes
To conclude this section, let us briefly discuss a somewhat related topic, the quasi-normal modes of black holes (and their analogues). In the discussion above, we explained super-radiance via the scattering picture where an incoming wave packet is scattered at the inhomogeneous flow profile (2.28) and thus splits (or is torn apart) into an in-falling and an outgoing part. Of course, using such a scattering picture, one is led to the question of whether there are bound states – which would then affect the dynamics. Even though there are typically no real bound states in such black-hole space times, there can be quasi-bound states or quasi-normal modes.
@@
In order to sketch the main idea by means of a simple picture, let us consider the one-dimensional Schrödinger equation in the presence of a rectangular (i.e., box) potential barrier such as $V(x)=V_{0}\Theta(L^{2}-x^{2})$. In classical mechanics, a particle with an energy $E>V_{0}$ starting at $x=0$ with $\dot{x}>0$ would just fall off the potential edge at $x=L$ and continue its propagation. In quantum mechanics, however, we have the phenomenon of quantum reflection and the particle can actually be reflected at the potential edge. Thus, we may approximate the two potential edges at $x=\pm L$ by (imperfectly) reflecting boundary conditions – where we find approximate bound state at quasi-discrete energy levels. However, since there is also a non-vanishing transmission probability at the edges, these are not exact bound state because the wave function would slowly leak out. They are quasi-bound states or quasi-normal modes or resonances, characterized by a discrete set of frequencies with real and imaginary part. These frequencies allow us to infer properties of the potential. Roughly speaking, the spacing between the real parts tells us the potential length $L$ while the lowest real part and the imaginary parts are related to the potential height $V_{0}$.
@@
Very analogous phenomena occur in other wave equations, such as that for a mass-less and minimally coupled scalar field (2.1) in the Schwarzschild metric, for example. In terms of the Regge-Wheeler tortoise coordinate $r_{*}$ and after the usual separation ansatz $\phi(t,r_{*},\vartheta,\varphi)=e^{-i\omega t}\phi_{\omega,\ell}(r_{*})Y_{\ell,m}(\vartheta,\varphi)$, the remaining equation for $\phi_{\omega,\ell}(r_{*})$ can be cast into a Schrödinger-like form with an effective potential $V_{\rm eff}(r_{*})$. In the case of gravitational waves $h_{\mu\nu}$, the underlying wave equation is a bit more complicated, but it also supports quasi-normal modes. Actually, they are related to the ring-down signal at late times which has been observed at LIGO for the black-hole merging event [182].
@@
Even though quasi-normal modes can already be discussed on the level of purely classical wave equations, it has been speculated that they might be related to black-hole quantization, see, e.g., [183]. For this specific relation, black-hole analogues can probably not be used as a fully quantitative analogue, because they do not reproduce the Einstein equations (which are presumably important for black-hole quantization). Nevertheless, black-hole analogues can serve as a qualitative analogue for quasi-normal modes themselves, see, e.g., [184, 185, 186, 187, 188].
@@ 2.8.1   Sauter-Schwinger effect
However, building a quantum simulator based on ultra-cold atoms for the dynamics in Eq. (2.30) may be non-trivial due to the second derivative terms. Thus, let us instead focus on the Dirac equation in 1+1 dimensions from Eq. (1.19) as already discussed in Sec. 1.2.4
@@
However, as a major distinction to ordinary tunneling in non-relativistic quantum mechanics, here the tunneling occurs between the negative continuum $\omega-q\phi<-m$ (i.e., the Dirac sea) and the positive continuum $\omega-q\phi>+m$, as in Klein tunneling. Thus, we arrive at the following picture: Even in the vacuum state, the Dirac sea is filled with electrons and due to the electric field $E$, it can happen that one of these electrons tunnels through the classically forbidden region (i.e., the mass gap between $-m$ and $+m$) into the positive continuum, where it emerges as a real particle. Since the total number of fermions is conserved, it leaves behind a hole in the Dirac sea – which is then a positron. Thus, the electric field $E$ should be able to spontaneously create electron-positron pairs out of the vacuum, as depicted in Fig. 2.5.
@@ 2.8.3   Quantum simulator for Sauter-Schwinger and Breit-Wheeler effect
As already explained in Sec. 1.2.4, the 1+1 Dirac Hamiltonian (1.19) can modeled by fermionic atoms in a bi-chromatic optical lattice, see Fig. 1.4. Quite intuitively, an electric field can then be simulated by applying a force to the atoms in the lattice – such that we arrive at a quantum simulator as sketched in Fig. 2.6, see also [75, 76]. The bosonic [^20] Due to the Pauli principle, bosonic atoms can be cooled down much better than fermionic atoms. analogue of such a simulator has been realized in another experiment [195] at Maryland (not to be confused with the expanding Universe simulator [150]). Since the bosonic (Rubidium) atoms used in this experiment [195] do not obey the Pauli principle, one does not have the precise analogue of the Dirac sea as sketched above. Therefore, the experiment [195] instead created an initial state where all the states in the lower band (which would correspond to the Dirac sea) were occupied by some bosons (with a certain probability) while the upper band was still empty. To ensure that the upper band was not excited in this initialization process, they started with a large gap, as in Fig. 2.6 (top left), and later adiabatically reduced it to the desired value, as in Fig. 2.6 (top right), thereby enabling tunneling transitions from the lower band to the upper band, i.e., the analogue of the Sauter-Schwinger effect [^21] As one might expect, tunneling from the valence band to the conduction band as an analogue of the Sauter-Schwinger effect is not restricted to ultra-cold atoms, similar effects can also occur in other systems, such as semi-conductors, see, e.g., [196] and references therein..
@@ 2.9   Dynamical Casimir effect
In order to visualize this attractive force, it is sometimes argued that there are less modes allowed in between the mirrors than outside – and that, therefore, the vacuum energy must be smaller. While it is correct that there are less allowed modes in between the mirrors (perpendicular to the mirrors, we have a discrete spectrum inside but a continuous spectrum outside), this picture is oversimplified: Even though there are less modes, they have a higher weight and thus the calculation of the vacuum energy requires a subtle subtraction procedure (where the infinite vacuum energy cancels). The failure of the above oversimplified arguments can be illustrated by comparing different boundary conditions at the mirrors. While two mirrors with the same boundary conditions (e.g., both Dirichlet or both Neumann) indeed attract each other, two different mirrors, one with Dirichlet and one with Neumann boundary conditions, repel each other. Similarly, two dielectric or two paramagnetic media typically attract each other while a dielectric and a paramagnetic medium tend to have a repulsive Casimir force.
@@
As a result, although the static Casimir effect has been verified experimentally (see, e.g., [207]), the situation for the dynamical Casimir effect is more complicated. The analogue of the dynamical Casimir effect has been observed in the Gothenburg [208] and Helsinki [209] experiments in time-dependent electromagnetic wave guides [^22] Electromagnetic wave guides with suitable properties could also be used to generate the analogue of Hawking radiation, see also [210]., but a conclusive verification of photon pair creation due to the real motion of a mirror in vacuum remains yet to be observed. As another point, one could argue that the well-known phenomenon of parametric down conversion in non-linear optical media can be interpreted as the creation of photon pairs by a spatio-temporal change of the refractive index $n(t,\mbox{\boldmath$r$})$.
@@ 3.2   Kibble-Zurek mechanism
The Kibble-Zurek mechanism [228, 229, 230, 231] refers to the creation of topological defects is symmetry-breaking phase transitions. Before turning to their creation mechanism, let us briefly discuss the concept of topological defects, where we focus on classical fields for the moment. Examples for topological defects are the kinks and anti-kinks discussed in the previous section, as well as vortices in Bose-Einstein condensates discussed in Sec. 2.7.1. Generally speaking, topological defects can occur in a phase of broken symmetry where the ground state is not unique but we have a set or manifold of ground states. In such a case, the vacuum would be a state with the same position in the ground-state manifold everywhere – often called the order parameter. A topological defect is then an inhomogeneous deviation from this homogeneous vacuum state where different spatial points assume distinct order parameters such that smooth deformations of the order parameter cannot transform the defect away (i.e., it is topologically protected).
As a rule of thumb [^23] A more rigorous approach can be based on homotopy groups or homotopy classes., topological defects can be characterized by the dimension $D_{\rm gs}$ of the ground-state manifold in the broken-symmetry phase and the space-time dimension $D_{\rm st}$, which then determine the dimensionality $D_{\rm td}$ of possible topological defects via $D_{\rm td}=D_{\rm st}-D_{\rm gs}-2$. Kinks and anti-kinks in the sine-Gordon model (3.1) correspond to a breaking of the discrete $\mathbb{Z}$ symmetry $\phi\to\phi+2\pi\mathbb{Z}$ such that ground-state manifold is discrete, i.e., zero-dimensional $D_{\rm gs}=0$. Thus, in 1+1 dimensions $D_{\rm st}=2$, they are point-like defects (i.e., also zero-dimensional) $D_{\rm td}=0$. In 2+1 dimensions $D_{\rm st}=3$, they would be line-like defects (i.e., one-dimensional) $D_{\rm td}=1$, and so on. As our second example, vortices in Bose-Einstein condensates correspond to a breaking of the $U(1)$ symmetry of the original Hamilton (1.4) by the phase of the condensate wave function $\psi_{\rm c}$. Thus, the ground-state manifold is one-dimensional, i.e., a circle, $D_{\rm gs}=1$. As already discussed in Sec. 2.7.1, a vortex is a point-like defect $D_{\rm td}=0$ in 2+1 dimensions $D_{\rm st}=3$ and a line-like defect $D_{\rm td}=1$ in 3+1 dimensions $D_{\rm st}=4$. As an intuitive picture, if we encircle a (singly quantized) vortex in real space, the phase (i.e., order parameter) describes a circle in order-parameter space. For multiply quantized vortices, the phase would wind around the circle in internal space several times, where the number of revolutions corresponds to the winding number, i.e., the topological charge.

### v2

@@ 2.4.1   Analogue Universe in condensate at rest
Instead of a monotonically expanding Universe, an oscillating scale factor has been simulated in the Chicago experiment [154] by modulating the external magnetic field near a Feshbach resonance [103]. By measuring the number of excited atoms in situ, it was possible to observe the thermal character of the radiation emitted into a certain direction. In terms of the coupled harmonic oscillator picture sketched in Fig. 2.1, the reduced density matrix of one oscillator is thermal, i.e., a mixed state, even though the total quantum state of the two oscillators is a pure (entangled) state. Indeed, the coherence associated to this two-mode squeezed state could also be tested in the Chicago experiment [154]. Note that the title of the publication [154] refers to Unruh radiation, which can probably be explained by the fact that squeezing underlies both phenomena, cosmological particle creation and the Unruh effect (as well as many others), but the analogy to cosmological particle creation as considered in this section is apparently closer and more quantitative.
@@ 2.4.2   Expanding condensates
An expanding ring-shaped Bose-Einstein condensate has been realized in the Maryland experiment [160, 161]. Since the wave-length of the phonons expands with the condensate, such a set-up is well suited for observing the red-shift of the field modes within an expanding Universe. In the Maryland experiment [160, 161], this has been achieved by imprinting a sound wave (i.e., a classical excitation) in the initial state of the condensate. As a related but more subtle phenomenon, the Maryland experiment [160, 161] did also find evidence for Hubble friction, see also [162]. To explain this phenomenon, let us consider the equation of motion (2.1) for a mass-less scalar field in the Friedmann-Lemaitre-Robertson-Walker metric (2.18) in 3+1 dimensions in terms of the proper time $\tau$
@@
After the initial period of linear evolution, the Maryland experiment [160, 161] was also able to observe the subsequent non-linear dynamics, which are qualitatively analogous to the pre-heating and re-heating periods in the early Universe after the end of inflation, see also [163].
@@ 2.5   Gibbons-Hawking effect
at the expense of introducing coordinate singularities at the de Sitter horizon $r=\pm 1/H$. In 3+1 dimensions, we would find the same form with an additional $r^{2}d\Omega^{2}$ for the angular part.
@@ 2.7   Super-radiance and Penrose process
So far, the phenomenon of super-radiance discussed above refers to purely classical wave dynamics. Actually, it even has an analogue in terms of classical particles, which is the Penrose process [171] and allows us to extract energy from a rotating black hole. Furthermore, the phenomenon of super-radiance can be generalized to other scenarios such as a rotating body [172]. Generally speaking, super-radiance requires an effective energy which can become negative, some sort of scattering or tearing apart process and some mechanism to absorb the parts with negative energy. For surface waves [170], this phenomenon has actually been observed for such a vortex flow (2.28) in the Nottingham experiment [173]. Replacing the water as used in the Nottingham experiment [173] by a “photon super-fluid”, an analogous effect has been observed in the Glasgow experiment [174], see also [175, 176, 177, 178] and [179] as well as [180].
@@ 2.7.2   Quasi-normal modes
To conclude this section, let us briefly discuss a somewhat related topic, the quasi-normal modes of black holes (and their analogues), see, e.g., [198, 199, 200]. In the discussion above, we explained super-radiance via the scattering picture where an incoming wave packet is scattered at the inhomogeneous flow profile (2.28) and thus splits (or is torn apart) into an in-falling and an outgoing part. Of course, using such a scattering picture, one is led to the question of whether there are bound states – which would then affect the dynamics. Even though there are typically no real bound states in such black-hole space times, there can be quasi-bound states or quasi-normal modes.
@@
In order to sketch the main idea by means of a simple picture, let us consider the one-dimensional Schrödinger equation in the presence of a rectangular (i.e., box) potential barrier such as $V(x)=V_{0}\Theta(L^{2}-x^{2})$. In classical mechanics, a particle with an energy $E>V_{0}$ starting at $x=0$ with $\dot{x}>0$ would just fall off the potential edge at $x=L$ and continue its propagation. In quantum mechanics, however, we have the phenomenon of quantum reflection and the particle can actually be reflected at the potential edge. Thus, we may approximate the two potential edges at $x=\pm L$ by (imperfectly) reflecting boundary conditions – where we find approximate bound state at quasi-discrete energy levels. However, since there is also a non-vanishing transmission probability at the edges, these are not exact bound state because the wave function would slowly leak out. They are quasi-bound states or quasi-normal modes or resonances, characterized by a discrete set of frequencies with real and imaginary parts. These frequencies allow us to infer properties of the potential. Roughly speaking, the spacing between the real parts tells us the potential length $L$ while the lowest real part and the imaginary parts are related to the potential height $V_{0}$.
@@
Very analogous phenomena occur in other wave equations, such as that for a mass-less and minimally coupled scalar field (2.1) in the Schwarzschild metric, for example. In terms of the Regge-Wheeler tortoise coordinate $r_{*}$ and after the usual separation ansatz $\phi(t,r_{*},\vartheta,\varphi)=e^{-i\omega t}\phi_{\omega,\ell}(r_{*})Y_{\ell,m}(\vartheta,\varphi)$, the remaining equation for $\phi_{\omega,\ell}(r_{*})$ can be cast into a Schrödinger-like form [^22] Note that one should be a bit careful when relating solutions (even for fixed $\omega$) of the Klein-Fock-Gordon equation (2.1), which is relativistic, to those of the Schrödinger equation, which is non-relativistic, especially if they are not exactly stationary since $\omega$ is not purely real. However, the main ideas do still apply. with an effective potential $V_{\rm eff}(r_{*})$. In the case of gravitational waves $h_{\mu\nu}$, the underlying wave equation is a bit more complicated, but it also supports quasi-normal modes. Actually, they are related [^23] In analogy to the spectroscopy of atoms and molecules, this relation is exploited in the field of “black-hole spectroscopy”, see, e.g., [201, 202]. to the ring-down signal at late times which has been observed at LIGO for the black-hole merging event [203].
@@
Even though quasi-normal modes can already be discussed on the level of purely classical wave equations, it has been speculated that they might be related to black-hole quantization, see, e.g., [204]. For this specific relation, black-hole analogues can probably not be used as a fully quantitative analogue, because they do not reproduce the Einstein equations (which are presumably important for black-hole quantization). Nevertheless, black-hole analogues can serve as a qualitative analogue for quasi-normal modes themselves, see, e.g., [205, 206, 207, 208, 209, 210].
@@ 2.8.1   Sauter-Schwinger effect
However, building a quantum simulator based on ultra-cold atoms for the dynamics in Eq. (2.30) may be non-trivial242424 For example, the non-relativistic (many-body) Schrödinger equation for the atoms contains a first-order time derivative while the relativistic Eq. (2.30) is of second order in time (and space). . Thus, let us instead focus on the Dirac equation in 1+1 dimensions from Eq. (1.19) as already discussed in Sec. 1.2.4
@@
However, as a major distinction to ordinary tunneling in non-relativistic quantum mechanics, here the tunneling occurs between the negative continuum $\omega-q\phi<-m$ (i.e., the Dirac sea) and the positive continuum $\omega-q\phi>+m$, as in Klein tunneling. Thus, we arrive at the following picture252525 It should be stressed here that the picture of the Dirac sea is a means to visualize quantum vacuum fluctuations, but one should be careful when taking it too seriously. For example, besides an infinite vacuum energy and mass, it would also correspond to an infinite vacuum charge. Furthermore, in view of the $CPT$ symmetry, one could also say that the Dirac sea is full of positrons and the electrons are the holes.: Even in the vacuum state, the Dirac sea is filled with electrons and due to the electric field $E$, it can happen that one of these electrons tunnels through the classically forbidden region (i.e., the mass gap between $-m$ and $+m$) into the positive continuum, where it emerges as a real particle. Since the total number of fermions is conserved, it leaves behind a hole in the Dirac sea – which is then a positron. Thus, the electric field $E$ should be able to spontaneously create electron-positron pairs out of the vacuum, as depicted in Fig. 2.5.
@@ 2.8.3   Quantum simulator for Sauter-Schwinger and Breit-Wheeler effect
As already explained in Sec. 1.2.4, the 1+1 Dirac Hamiltonian (1.19) can modeled by fermionic atoms in a bi-chromatic optical lattice, see Fig. 1.4. Quite intuitively, an electric field can then be simulated by applying a force to the atoms in the lattice – such that we arrive at a quantum simulator as sketched in Fig. 2.6, see also [75, 76]. The bosonic [^29] Due to the Pauli principle, bosonic atoms can be cooled down much better than fermionic atoms. analogue of such a simulator has been realized in another experiment [217] at Maryland (not to be confused with the expanding Universe simulator [160, 161]). Since the bosonic (Rubidium) atoms used in this experiment [217] do not obey the Pauli principle, one does not have the precise analogue of the Dirac sea as sketched above. Therefore, the experiment [217] instead created an initial state where all the states in the lower band (which would correspond to the Dirac sea) were occupied by some bosons (with a certain probability) while the upper band was still empty. To ensure that the upper band was not excited in this initialization process, they started with a large gap, as in Fig. 2.6 (top left), and later adiabatically reduced it to the desired value, as in Fig. 2.6 (top right), thereby enabling tunneling transitions from the lower band to the upper band, i.e., the analogue of the Sauter-Schwinger effect [^30] As one might expect, tunneling from the valence band to the conduction band as an analogue of the Sauter-Schwinger effect is not restricted to ultra-cold atoms, similar effects can also occur in other systems, such as semi-conductors, see, e.g., [218] and references therein..
@@ 2.9   Dynamical Casimir effect
In order to visualize this attractive force, it is sometimes argued that there are less modes allowed in between the mirrors than outside – and that, therefore, the vacuum energy must be smaller. While it is correct that there are fewer allowed modes in between the mirrors (perpendicular to the mirrors, we have a discrete spectrum inside but a continuous spectrum outside), this picture is oversimplified: Even though there are fewer modes, they have a higher weight and thus the calculation of the vacuum energy requires a subtle subtraction procedure (where the infinite vacuum energy cancels). The failure of the above oversimplified arguments can be illustrated by comparing different boundary conditions at the mirrors. While two mirrors with the same boundary conditions (e.g., both Dirichlet or both Neumann for the scalar field case) indeed attract each other, two different mirrors, one with Dirichlet and one with Neumann boundary conditions, repel each other, see. e.g., [228]. Similarly, two dielectric or two paramagnetic media typically attract each other while a dielectric and a paramagnetic medium tend to have a repulsive Casimir force, cf. [229].
@@
As a result, although the static Casimir effect has been verified experimentally (see, e.g., [231]), the situation for the dynamical Casimir effect is more complicated. The analogue of the dynamical Casimir effect has been observed in the Gothenburg [232] and Helsinki [233] experiments in time-dependent electromagnetic wave guides [^31] Electromagnetic wave guides with suitable properties could also be used to generate the analogue of Hawking radiation, see also [234]., but a conclusive verification of photon pair creation due to the real motion of a mirror in vacuum remains yet to be observed. As another point, one could argue that the well-known phenomenon of parametric down conversion in non-linear optical media [^32] Note that the mapping of such non-linear media to refractive indices $n$ is very simplified: Depending on the symmetries of the medium, one should distinguish between $\chi^{(2)}$ and $\chi^{(3)}$ media, for example, where the material properties are typically encoded in tensorial quantities such as $\chi^{(2)}_{ijk}$ and $\chi^{(3)}_{ijkl}$ etc. can be interpreted as the creation of photon pairs by a spatio-temporal change of the refractive index $n(t,\mbox{\boldmath$r$})$.
@@ 3.2   Kibble-Zurek mechanism
The Kibble-Zurek mechanism [252, 253, 254, 255] refers to the creation of topological defects in symmetry-breaking phase transitions. Before turning to their creation mechanism, let us briefly discuss the concept of topological defects, where we focus on classical fields for the moment. Examples for topological defects are the kinks and anti-kinks discussed in the previous section, as well as vortices in Bose-Einstein condensates discussed in Sec. 2.7.1. Generally speaking, topological defects can occur in a phase of broken symmetry where the ground state is not unique but we have a set or manifold of ground states. In such a case, the vacuum would be a state with the same position in the ground-state manifold everywhere – often called the order parameter. A topological defect is then an inhomogeneous deviation from this homogeneous vacuum state where different spatial points assume distinct order parameters such that smooth deformations of the order parameter cannot transform the defect away (i.e., it is topologically protected).
As a rule of thumb [^33] A more rigorous approach can be based on homotopy groups or homotopy classes., topological defects can be characterized by the dimension $D_{\rm gs}$ of the ground-state manifold in the broken-symmetry phase and the space-time dimension $D_{\rm st}$, which then determine the dimensionality $D_{\rm td}$ of possible topological defects via $D_{\rm td}=D_{\rm st}-D_{\rm gs}-2$. Kinks and anti-kinks in the sine-Gordon model (3.1) correspond to a breaking of the discrete $\mathbb{Z}$ symmetry $\phi\to\phi+2\pi\mathbb{Z}$ such that ground-state manifold is discrete, i.e., zero-dimensional $D_{\rm gs}=0$. Thus, in 1+1 dimensions $D_{\rm st}=2$, they are point-like defects (i.e., also zero-dimensional) $D_{\rm td}=0$. In 2+1 dimensions $D_{\rm st}=3$, they would be line-like defects (i.e., one-dimensional) $D_{\rm td}=1$, and so on. As our second example, vortices in Bose-Einstein condensates correspond to a breaking of the $U(1)$ symmetry of the original Hamiltonian (1.4) by the phase of the condensate wave function $\psi_{\rm c}$. Thus, the ground-state manifold is one-dimensional, i.e., a circle, $D_{\rm gs}=1$. As already discussed in Sec. 2.7.1, a vortex is a point-like defect $D_{\rm td}=0$ in 2+1 dimensions $D_{\rm st}=3$ and a line-like defect $D_{\rm td}=1$ in 3+1 dimensions $D_{\rm st}=4$. As an intuitive picture, if we encircle a (singly quantized) vortex in real space, the phase (i.e., order parameter) describes a circle in order-parameter space. For multiply quantized vortices, the phase would wind around the circle in internal space several times, where the number of revolutions corresponds to the winding number, i.e., the topological charge.